# Testing theories of the glass transition with the same liquid, but many kinetic rules

### 22 Feb 2024 v3

@@ Abstract
Understanding why liquids glass formers cease to flow near their glass transition $T_{g}$ remains a challenge. At that point, the relaxation time $\tau_{\alpha}$ beyond which stress relaxes is of order of minutes, which is fifteen decades larger than at high temperatures. From $\tau_{\alpha}$, the activation energy $E_{a}$ can be defined as $\tau_{\alpha}=t_{0}\exp(E_{a}/T)$, where $t_{0}$ is a microscopic time scale and $T$ is the temperature (in the units of the Boltzmann constant). In liquids called fragile, $E_{a}$ can increase five-fold or more under cooling Anderson (1995); Ediger et al. (1996); Debenedetti and Stillinger (2001); Berthier and Biroli (2011). As the dynamics slows down, it also becomes more and more heterogeneous, corresponding to a growing length scale $\xi$  Kob et al. (1997); Yamamoto and Onuki (1998); Dalle-Ferrier et al. (2007); Karmakar et al. (2014). Contrasting theories seek to explain these two facts. In the first class of views, including Adam-Gibbs Adam and Gibbs (1965) and Random First Order Theory (RFOT) Kirkpatrick et al. (1989); Lubchenko and Wolynes (2007); Biroli and Bouchaud (2012), the increase of activation energy stems from the emergence of some order on a growing length $\xi$, that must be destroyed by cooperative motion on that scale to relax the material. $E_{a}$ can then be expressed in terms of purely thermodynamic quantities, independently of the kinetic rules governing the dynamics. Some real space approaches associate such a growing order to locally favored structures Patrick Royall et al. (2008); Kawasaki et al. (2007). A second viewpoint seeks to capture the mechanism of dynamical facilitation, whereby the relaxation of a given region speeds up the relaxation of regions nearby. Kinetically constrained models Ritort and Sollich (2003); Berthier et al. (2011), such as the East model, capture this effect and suggest a scenario Garrahan and Chandler (2002); Garrahan et al. (2007); Hedges et al. (2009) in which thermodynamics plays almost no role, but dynamics is heterogeneous and the growth of activation energy stems from non-local rearrangements taking place over $\xi$. At odd with these two views, a third approach that includes free volume Turnbull and Cohen (1961) or elastic Dyre (2006); Kapteijns et al. (2021); Lerbinger et al. (2022); Ciamarra et al. (2023) models, assumes that the activation energy is not controlled by a growing length scale. Instead, it is governed by the energy cost of elementary rearrangements of a few particles jumping over a barrier. The elastic coupling between rearrangements Lemaître (2014); Chowdhury et al. (2016); Tong et al. (2020); Wu et al. (2015); Maier et al. (2017); Steffen et al. (2022); Flenner and Szamel (2015); Klochko et al. (2022); Chacko et al. (2021) leads to a correlated dynamics  Ozawa and Biroli (2023) that can be described in terms of avalanches of activated events Tahaei et al. (2023).
@@
Figure 2: $\tau_{\alpha}/\tau_{\alpha}^{*}$ as function of $\phi$ for different values of $f_{R}$ and $f_{P}$ for $d=2$ (a) and $d=3$ (e). From these curves, $\phi_{G}$ as a function of $f_{R}$ and $f_{P}$ is extracted, as indicated in color for $d=2$ (b) and $d=3$ (f). The discs mark the simulation data, from which extrapolations are made. (c,g) show $\phi_{G}$ as a function of the number of degrees of freedom per particle $d+1-f_{R}-f_{P}$ and (d,h) show that $\phi_{G}$ can be collapsed as a function of an effective number of constraints $x=f_{P}+C(d)f_{P}$, with $C(2)=0.45$, and $C(3)=0.70$
@@
where $\xi$ is the correlation length, $\tilde{\sigma}$ some exponent and $\tilde{d}_{f}$ is the fractal dimension of the clusters. In this view, $\xi^{\tilde{d}_{f}}\sim\chi_{4}^{*}$ which characterizes the typical cluster volume. For $d=2$, $\tilde{d}_{f}\approx 2$ and $\tilde{\sigma}\approx 1/2$ Tahaei et al. (2023), leading to $l_{c}\sim\sqrt{\chi_{4}^{*}}\left[\ln(\tau_{\alpha}/t)\right]^{-1}$. As shown in in Fig.3(d), this prediction works remarkably well, independently of $f_{R}$. It supports that thermal avalanches cause DH independently of the kinetic rules chosen. Discussion: Swap Monte Carlo algorithms can be restricted to local moves with no significant effects on the dynamics Ninarello et al. (2017). More fundamentally, they are equivalent to a purely local kinetic rule where particles can adapt their radii Briano and Glandt (1984); Brito et al. (2018). For theories based on a growing order on some length $\xi_{coop}$ such as RFOT, or approaches based on locally favored structures, barriers are cooperative and can be expressed in terms of thermodynamic quantities alone. They should be present for any local kinetic rules Wyart and Cates (2017), including those studied here. Thus in these approaches, the core mechanism slowing down the dynamics near the glass transition should not depend on the choice of $(f_{R},f_{P})$. Authors in Berthier et al. (2019) acknowledged that swap and normal dynamics should asymptotically relax at the same pace according to RFOT, but countered that pre-asymptotic corrections (not currently described within this theory) may cause the observed difference. The main difficulty with this view is that the dynamics is very different as the parameters $(f_{R},f_{P})$ change: as shown in Fig.5, these dynamics do not become equivalent even after a slowing down of 15 decades accessible experimentally. To have predictive power, RFOT or theories based on thermodynamic quantities should specify a value for the parameters $(f_{R},f_{P})$ for which they apply. However, currently they don’t. Our observation that $\phi_{G}$ continuously depends on $(f_{R},f_{P})$, and does not plateau to some constant value around the normal dynamics $(1,0)$, shows that the normal dynamics is just one among many other dynamics. This point underlines the lack of predictive power of RFOT or related theories - at least for the systems of continuously polydisperse particles studied here. By contrast, for theories based on kinetic constraints or on local barriers (known to depend on kinetic rules Brito et al. (2018)), the fact that $\phi_{G}$ should continuously vary with the amount of constraints is evident. The normal dynamics is slower simply because it is a kinetically constrained version of swap dynamics. Conclusion: RFOT is a mean-field theory of the glass transition, which has shown undeniable successes. It is exact in infinite dimension Charbonneau et al. (2014); Maimbourg et al. (2016), correctly captures aspects of the thermodynamics of super-cooled liquids Biroli et al. (2008), and presents a dynamical transition Kurchan and Laloux (1996); Biroli et al. (2006); Lubchenko and Wolynes (2007) akin to mode coupling theory, that describes some aspects of liquid dynamics at intermediate temperatures Götze (2008). Yet our results support that its description of activation near the glass transition does not apply for the continuously polydisperse systems studied here. Although our conclusions are restricted to these specific systems where swap is so performant, these models are known to capture the key facts associated with the glass transition Scalliet et al. (2022). The model we introduced, with its very large variation of dynamics with different kinetic rules, offers new opportunities to test different theories of the glass transition, extending previous observations that were only considered for a single kinetic rule. As an illustration, we showed that the time evolution of dynamical heterogeneities follow a description of facilitation based on thermal avalanches of local rearrangements Tahaei et al. (2023) independently of the kinetic rules chosen, emphasizing the robustness of this theory. Other debated issues that this model would help resolve include what observable predicts best the regions which will flow first, the so-called dynamical propensity, for which many candidates were proposed Widmer-Cooper et al. (2004); Lerbinger et al. (2022); Alkemade et al. (2022); Schoenholz et al. (2017). Likewise, numerical tests put forward to test theories of the slowing down of the dynamics (i.e. how the activation energy $E_{a}$ depends on temperature or packing fraction) can now be made much more stringent. The list includes kinetically constrained or lattice gas models Toninelli et al. (2006), that can be tested by analyzing irreversible events Keys et al. (2011). Furthermore, the notion that glassy dynamics correspond to local rearrangements was supported by the measurement of the density of state of local barriers Ciamarra et al. (2023); and these barriers were argued to be governed alternatively by global Dyre (2006) or local Kapteijns et al. (2021) elasticity, or by the amplitude of vibrational motion Puosi and Leporini (2012). Varying $(f_{R},f_{P})$ in these measurements will indicate which viewpoint is most likely correct. Overall, we have added an axis to the glass transition problem by varying continuously kinetic rules, affecting strongly observations and giving a new handle to decide which theory of the glass transition actually applies.

### 27 Feb 2024 v4

@@ Abstract
Understanding why liquids glass formers cease to flow near their glass transition $T_{g}$ remains a challenge. At that point, the relaxation time $\tau_{\alpha}$ beyond which stress relaxes is of order of minutes, which is fifteen decades larger than at high temperatures. From $\tau_{\alpha}$, the activation energy $E_{a}$ can be defined as $\tau_{\alpha}=t_{0}\exp(E_{a}/T)$, where $t_{0}$ is a microscopic time scale and $T$ is the temperature (in the units of the Boltzmann constant). In liquids called fragile, $E_{a}$ can increase five-fold or more under cooling Anderson (1995); Ediger et al. (1996); Debenedetti and Stillinger (2001); Berthier and Biroli (2011). As the dynamics slows down, it also becomes more and more heterogeneous, corresponding to a growing length scale $\xi$  Kob et al. (1997); Yamamoto and Onuki (1998); Dalle-Ferrier et al. (2007); Karmakar et al. (2014). Contrasting theories seek to explain these two facts. In the first class of views, including Adam-Gibbs Adam and Gibbs (1965) and Random First Order Theory (RFOT) Kirkpatrick et al. (1989); Lubchenko and Wolynes (2007); Biroli and Bouchaud (2012), the increase of activation energy stems from the emergence of some order on a growing length $\xi$, that must be destroyed by cooperative motion on that scale to relax the material. $E_{a}$ can then be expressed in terms of purely thermodynamic quantities, independently of the kinetic rules governing the dynamics. Some real space approaches associate such a growing order to locally favored structures Patrick Royall et al. (2008); Kawasaki et al. (2007). A second viewpoint seeks to capture the mechanism of dynamical facilitation, whereby the relaxation of a given region speeds up the relaxation of regions nearby. Kinetically constrained models Ritort and Sollich (2003); Berthier et al. (2011), such as the East model, capture this effect and suggest a scenario Garrahan and Chandler (2002); Garrahan et al. (2007); Hedges et al. (2009) in which thermodynamics plays almost no role, but dynamics is heterogeneous and the growth of activation energy stems from non-local rearrangements taking place over $\xi$. At odds with these two views, a third approach that includes free volume Turnbull and Cohen (1961) or elastic Dyre (2006); Kapteijns et al. (2021); Lerbinger et al. (2022); Ciamarra et al. (2023) models, assumes that the activation energy is not controlled by a growing length scale. Instead, it is governed by the energy cost of elementary rearrangements of a few particles jumping over a barrier. The elastic coupling between rearrangements Lemaître (2014); Chowdhury et al. (2016); Tong et al. (2020); Wu et al. (2015); Maier et al. (2017); Steffen et al. (2022); Flenner and Szamel (2015); Klochko et al. (2022); Chacko et al. (2021) leads to a correlated dynamics  Ozawa and Biroli (2023) that can be described in terms of avalanches of activated events Tahaei et al. (2023).
@@
Figure 2: $\tau_{\alpha}/\tau_{\alpha}^{*}$ as function of $\phi$ for different values of $f_{R}$ and $f_{P}$ for $d=2$ (a) and $d=3$ (e). From these curves, $\phi_{G}$ as a function of $f_{R}$ and $f_{P}$ is extracted, as indicated in color for $d=2$ (b) and $d=3$ (f). The discs mark the simulation data, from which extrapolations are made. (c,g) show $\phi_{G}$ as a function of the number of degrees of freedom per particle $d+1-f_{R}-f_{P}$ and (d,h) show that $\phi_{G}$ can be collapsed as a function of an effective number of constraints $x=f_{P}+C(d)f_{R}$, with $C(2)=0.45$, and $C(3)=0.70$. The normal dynamics is marked as a red square in these plots.
@@
where $\xi$ is the correlation length, $\tilde{\sigma}$ some exponent and $\tilde{d}_{f}$ is the fractal dimension of the clusters. In this view, $\xi^{\tilde{d}_{f}}\sim\chi_{4}^{*}$ which characterizes the typical cluster volume. For $d=2$, $\tilde{d}_{f}\approx 2$ and $\tilde{\sigma}\approx 1/2$ Tahaei et al. (2023), leading to $l_{c}\sim\sqrt{\chi_{4}^{*}}\left[\ln(\tau_{\alpha}/t)\right]^{-1}$. As shown in Fig.3(d), this prediction works remarkably well, independently of $f_{R}$. It supports that thermal avalanches cause DH independently of the kinetic rules chosen. Discussion: Swap Monte Carlo algorithms can be restricted to local moves with no significant effects on the dynamics Ninarello et al. (2017). More fundamentally, they are equivalent to a purely local kinetic rule where particles can adapt their radii Briano and Glandt (1984); Brito et al. (2018). For theories based on a growing order on some length $\xi_{coop}$ such as RFOT, or approaches based on locally favored structures, barriers are cooperative and can be expressed in terms of thermodynamic quantities alone. They should be present for any local kinetic rules Wyart and Cates (2017), including those studied here. Thus in these approaches, the core mechanism slowing down the dynamics near the glass transition should not depend on the choice of $(f_{R},f_{P})$. Authors in Berthier et al. (2019) acknowledged that swap and normal dynamics should asymptotically relax at the same pace according to RFOT, but countered that pre-asymptotic corrections (not currently described within this theory) may cause the observed difference. The main difficulty with this view is that the dynamics is very different as the parameters $(f_{R},f_{P})$ change: as shown in Fig.5, these dynamics do not become equivalent even after a slowing down of 15 decades accessible experimentally. To have predictive power, RFOT or theories based on thermodynamic quantities should specify a value for the parameters $(f_{R},f_{P})$ for which they apply. However, currently they don’t. Our observation that $\phi_{G}$ continuously depends on $(f_{R},f_{P})$, and does not plateau to some constant value around the normal dynamics $(1,0)$, shows that the normal dynamics is just one among many other dynamics. This point underlines the lack of predictive power of RFOT or related theories - at least for the systems of continuously polydisperse particles studied here. By contrast, for theories based on kinetic constraints or on local barriers (known to depend on kinetic rules Brito et al. (2018)), the fact that $\phi_{G}$ should continuously vary with the amount of constraints is evident. The normal dynamics is slower simply because it is a kinetically constrained version of swap dynamics. Conclusion: RFOT is a mean-field theory of the glass transition, which has shown undeniable successes. It is exact in infinite dimension Charbonneau et al. (2014); Maimbourg et al. (2016), correctly captures aspects of the thermodynamics of super-cooled liquids Biroli et al. (2008), and presents a dynamical transition Kurchan and Laloux (1996); Biroli et al. (2006); Lubchenko and Wolynes (2007) akin to mode coupling theory, that describes some aspects of liquid dynamics at intermediate temperatures Götze (2008). Yet our results support that its description of activation near the glass transition does not apply for the continuously polydisperse systems studied here. Although our conclusions are restricted to these specific systems where swap is so performant, these models are known to capture the key facts associated with the glass transition Scalliet et al. (2022). The model we introduced, with its very large variation of dynamics with different kinetic rules, offers new opportunities to test different theories of the glass transition, extending previous observations that were only considered for a single kinetic rule. As an illustration, we showed that the time evolution of dynamical heterogeneities follow a description of facilitation based on thermal avalanches of local rearrangements Tahaei et al. (2023) independently of the kinetic rules chosen, emphasizing the robustness of this theory. Other debated issues that this model would help resolve include what observable predicts best the regions which will flow first, the so-called dynamical propensity, for which many candidates were proposed Widmer-Cooper et al. (2004); Lerbinger et al. (2022); Alkemade et al. (2022); Schoenholz et al. (2017). Likewise, numerical tests put forward to test theories of the slowing down of the dynamics (i.e. how the activation energy $E_{a}$ depends on temperature or packing fraction) can now be made much more stringent. The list includes kinetically constrained or lattice gas models Toninelli et al. (2006), that can be tested by analyzing irreversible events Keys et al. (2011). Furthermore, the notion that glassy dynamics correspond to local rearrangements was supported by the measurement of the density of state of local barriers Ciamarra et al. (2023); and these barriers were argued to be governed alternatively by global Dyre (2006) or local Kapteijns et al. (2021) elasticity, or by the amplitude of vibrational motion Puosi and Leporini (2012). Varying $(f_{R},f_{P})$ in these measurements will indicate which viewpoint is most likely correct. Overall, we have added an axis to the glass transition problem by varying continuously kinetic rules, affecting strongly observations and giving a new handle to decide which theory of the glass transition actually applies.